# Does quantum information require additional structure?

### v2

@@ 5 Quantum model as a paradigm for quantum information
In contrast to an “unspeakable” quantum information, its mathematical representation $\psi$ can be broadcasted. A paradigmatic example is quantum cryptography considered in [82]. Namely, the key distribution schemes show that there is a feature that is naturally ascribed to the quantum information carriers (QIC). It is clear that there must be quantum communication between Alice and Bob to achieve secret key. Indeed, Alice cannot simply send a wave function $\psi$ written on paper to Bob, as Eve could read it out without disturbing it. However, by sending QIC, Alice and Bob can achieve something impossible within the “classical” world. Another example is the direct calculation of quantum state function $f(\varrho)$ [82]. It can be viewed as a kind of quantum computing on purely quantum input. This approach is natural if the quantum state $\varrho$ is viewed as representing real quantity to be processed rather than state of our knowledge.
The above examples show that the quantum information exists physically, but it cannot be identified with its mathematical representation $\psi$. Here, a natural question arises: How does this affect the standard understanding of the Born algorithm? It is not excluded that in contrast to the classical models in which the concept of the classical information and probability are on the same footing in the sense that they can be derived from one another. At the quantum level, the concept of probability has its formal root in Gleason’s theorem, while the notion of quantum information is still not well defined. The reason is that in the quantum models of closed systems, there is no general concept of joint and conditional probability cannot be defined for the general case except for the commuting observables [78]. Indeed, in the RC model a probabilistic measure defined on phase space differs radically from the quantum probabilistic measure. In fact, the latter is largely determined by the projection structure of the von Neumann algebra, and as it follows from Gleason’s theorem, therefore it cannot be freely defined. Quantum observables are not random variables in classical probabilistic space; there are no “elements of reality.” However, for the classical measure, there is great freedom in the way it is applied from the outside to the structure of the space of phase states. It seems that the “Born’s noncontextual probability” is not an autonomic notion but depends on nonclassical algebraic properties of quantum observables.
6 Nonrelativistic quantum model: laboratory space versus configuration space
@@
It is still remains unclear why the non-relativistic standard model already predicts the existence of stable quantum structures and EPR correlations and at the same time is consistent with RQF. Note that to protect the later from superluminal speeds, a priory postulate of local commutativity is necessary [83]. However, the RQF is not strictly limited by the speed of light [84]. In this context, the existence of the EPR correlations naturally lead to the question of their physical nature and still is subject to a wide debate [85]. In particular, Gisin suggested “that quantum correlations somehow arise from outside space-time”[86]. Clauser and Shimony had previously suggested that “Either we must completely abandon the realistic philosophy of almost all practicing physicists or fundamentally revise our concept of space-time” [87]. The second part of the above alternative is big challenge since Lorenz’s symmetry $M_{4}$ imposed by heave classical devices is an a priori measurement condition of any physical model (Bohr). On the other hand this challenge encounters an intriguing problem of incompatibility formulations of NRQ in laboratory space and configuration space, which has recently been the subject of deep consideration in the context of local realism and correlation EPR [88].
@@
Namely, according to Born’s original formulation, the wave function $\psi_{lab}(\mathbf{r}_{lab},t)$ represents a field that propagates in lab space and correspondingly the real-valued probability density $|\psi_{lab}(\mathbf{r}_{lab},t)|^{2}$ also propagates as a classical field. Thus, Born’s modification of the interpretation of the wave function is expressed only in its statistical interpretation. This is what limits lab-space formulation the most is that it only applies to the single particle systems, and there is no strict method to extend this formulation to $n\geq 2$ particle systems. In this context, Clauser notes “This impossibility might be expected since Local Realism’s formulation and Bell’s Theorem forbids it. Correspondingly, experiments that refute Local Realism also refute a lab space formulation of quantum mechanics.”
@@
Von Neumann introduced the abstract configuration-space formulation of quantum mechanics for the system consisting of $N$ particles, which uses a very general abstract $k$-dimensional argument space for the wave function that is called configuration space ($k=3N$), which enables a system with $k$-degrees of freedom: $\mathbf{r}_{1,conf},\ldots,\mathbf{r}_{N,conf}$, and this was generalized by incorporating $k$-quantum degree of freedom of the system: [89] $s_{1,conf},\ldots,s_{N,conf}$. Then, a configuration-space wave function can be written as: $\psi_{conf,N}(\mathbf{r}_{1,conf},\ldots,\mathbf{r}_{N,conf},s_{1,conf},\ldots,s_{N,conf})$. As Clauser notes, Born’s Lab-space formulation can only describe single-particle systems and it is essentially a form of BCHS local realism that is violated by quantum mechanics, while a completely abstract configuration space approach grasps the holistic phenomenon of entanglement, which is not a one-body phenomenon but always occurs in multipartite systems.
The inconsistency between the formulation of the laboratory space and the concept of configuration space, a.k.a Born’s argument space ambiguity shows implicitly that the fundamental dispute between Einstein and Bohr regarding the interpretation of the wave function was in fact, a dispute between Einstein, the “eventist” (“Alles ist Koinzidenz”) and Bohr, the “epistemic relationist”. Einstein believed that since measurement is subject to Lorentz $M_{4}$ symmetry, hence the classical “esse” coincides with “percipi” (measurement) in the laboratory space and this view does not contradict Born’s Lab-space statistical interpretation. In this context, it may be surprising that both model RQ and NRQ maintain the classical space-time continuum of directly observable events as an elementary background in the most fundamental models.
Whereas Bohr does not associate any reality with the wave function and treats it as intersubjective information about what can be updated in the measurement. Bohr $s$ view seems to be consistent with the configuration-space formulated wave function, as the latter does not provide a space-time description of the global wave function. Clauser concludes “if we assume, according to the Copenhagen interpretation of quantum mechanics, that abstract configuration space is inherently impenetrable, then it allows for magical action at a distance, and no objects are necessary to exist as things in the laboratory space” [88].
@@ 7 Quantum relational space hypothesis
Note, however, that if we accept the correspondence principle, we can further claim that the wave function $\psi_{conf}$ is a mathematical representation (image) of quantum information. Consequently, it must have a non-spatial nature. Unfortunately the $\psi_{conf}$ space formalism does not provide any conceptual model to understand its “inner workings”, although intuitively, it seems to hide a deeper level of reality. Despite the vast knowledge about entanglement, it is not still clear where it has its habitat and how it “coexists” with fundamental Galilean and Lorentz-Poincare symmetries. In 1999, Gisin’s team conducted a Michelson-Morely-like experiment which showed that the 2-photon interferences were still visible, independently of the relative velocity between Alice’s and Bob’s reference frames (actually, the magnitude of the velocity was constant at $100m/s$, but its orientation varied) [90] This result suggests that the von Neumann’s collapse of the quantized electromagnetic field cannot be viewed as a real causal process, and quantization of the field definitely renders it no longer describable as “real stuff in real space–time” [90, 88].
@@
The above conclusions incline us to postulate that the fundamental physical interactions between quantum objects generate directly unobservable relations, being a source of their physical properties. In particular, we assume that every completely isolated system can be associated with an intrinsic quantum space $Q_{rel}$ of directly unobservable relations generated by unitary interactions which generate a broad class of relations in which no subsystems are required to be macroscopic directly accessible to active observer. We assume that the space of relations $Q_{rel}$ involves intrinsic quantum degrees of freedom such as spin, color, and smell, and it precedes the space of directly observable correlations, $Q_{cor}$, which is some subspace of the “classical” Minkowski $M_{4}$ space-time (Fig.1).
@@
Figure 1: $Q_{rel}$ relation space of reversible “pre-measurements” precedes the space of correlation events $Q_{cor}$. Arrows indicate irreversible decoherence
We are not able to determine the mathematical structure of the hypothetical relational space $Q_{rel}$. The simplest way to get around this difficulty is to assume that the relational space structure is isomorphic to the Hilbert space tensorial structure. Then, we can say that quantum models provide instructions on how to model a structure of $Q_{rel}$ in terms of Hilbert space formalism. For instance, the algorithmic generation of genuine multipartite entangled states via indistinguishability of identical particles by coherent unitary deformation operations can be viewed as the kind of modeling of the $Q_{rel}$ space [91]. In particular, pure quantum state $\psi$ of the compound system encodes information about all relations (amplitude and phase) between its subsystems, which may be related to “the correlations without correlata” [9]. Obviously, the assumption on the ideal isolation of the system is realistic idealization, as the unitary interactions do not yield outcomes [92] and the system cannot “recognize” itself. It is intimately connected with the measurement problem – “Achilles’ heel” of quantum modeling of the world [28, 29, 30]. It should be noted here that for additive conserved quantities, the measurement problem is, in principle, solvable within ”Stinesprig” unitary representation based on dilated Hilbert space that includes both the degrees of freedom of the specimen and apparatus [93, 94]. In this approach, irreversible collapse does not appear from the Schrödinger equation but is induced then temporal evolution is constrained by conservation laws. In realistic scenarios, however, the observer is not able to control all the degrees of freedom environment of the system and broadcast of the results [95]. Thus, he is forced to use a formalism that includes open systems and irreversible decoherence (i.e., [96, 97]). Some authors assume [8, 98] that any physical system can play the role of an observer in a physical interaction that every interaction is a measurement regardless of the strength and time of the interaction between the systems and Bohr’s heavy basis conditions [99], which says that the apparatus is capable of realizing ideal space-time realization only if it is infinitely inert (See in this context [100, 101]).
@@
Here, we distinguish a class of projective, strong, and instantaneous measurements that satisfy the condition of heavy basis, which removes the quantum Doppler effect. This leads to a boundary factual-like type of correlations between a ”specimen” and measuring device where the latter is a subsystem large enough that an observer can perceive it directly. It is compatible with our assumption that the relation space $Q_{rel}$ precedes subspace $Q_{cor}$ of Minkowski space $M_{4}$ of the actualizations (clicks) appearing as a result of irreversible measurements with specially prepared macro-objects (detectors) in finite regions $C$ of $M_{4}$. Then, it is reasonable to introduce measure $\mu(C)$ as the probability of a ‘click’ provided that a suitable detector operates in region C [102]. This seems consistent with the suggestion [103] that the quantum probabilistic structure is “overbuilt” over $M_{4}$, while the latter is only a “carrier” for the probabilistic measures. Note, however, that the Hilbertian “isomorphic constraint” imposed on the structure of $Q{rel}$ may be too restrictive, which may rule out the description of potential “exotic” effects such as quantum gravity.
Finally, it is worth noting, that the hypothesis of relational space confirms Gisin’s conjecture that the “quantum correlations arise from outside space-time” and that idea of quantum relational space is compatible with Reichenbach common cause principle [104].
@@ 8 Chyliński’s quantum relational continuum space
In this section, we present Chyliński’s model of a relational continuum space based on the configuration space of relative coordinates. The NRQ model ($\hbar\neq 0$) maintains the classical continuum of directly observable events $X^{\mu}=(\mathbf{X},X_{0})$, $\mu=1,2,3,4$, as an elementary background of most physical models. On the other hand, according to the uncertainty relations, the exact determination of 4-momentum $P^{\mu}=(\mathbf{P},P_{0})$, $P^{\mu}(\Delta P^{c}\to 0)$ rules out the $x$ localization of its quantum carrier in $M_{4}$ since $\Delta X^{\mu}\sim h/\Delta P^{\mu}\to\infty$. It destroys in an uncontrollable way the “quantum-potential” structure $\psi$ of the isolated system under measurement. As one knows, the direct p-measurements of the scattered particles in the asymptotic region do not disturb the structure of $\psi$ state, which is consistent with the state of isolation of the investigated system. The typical momentum measurements require measuring position and inferring momentum, which, however, do not violate the uncertainty principle. Note that there are also momentum measurements that do not rely on measuring position based on the Doppler effect, the energy-sensitive single-photon detector[105].
It is not obvious that $L$-symmetry of “heavy bases” valid micro-states when it is extrapolated to the interaction region. As we know the quantum $p$-$x$ duality establishes one-to-one correspondence between the p and x representations of a Hilbert vector $|G\rangle$ as
@@
Note, however, that in this abstract formulation of the state $G$ of system $S_{M}$ the physical meaning of the $x$ and $p$ variables remain undetermined. According to the $M_{4}$ framework, the $p$’s denote four-momenta of the corresponding degree of freedom of $S_{M}$ represented in some reference frame $S$. However, 4-symmetry revealed the experimental role of the asymptotic language $p$ of relativistic kinematics, which was enhanced by relativistic S matric formalism. Thus, we can first measure the $S$ matrix elements (cross-sections) parametrized by Mandelsztam $M$-invariant variable $s_{k}$ determined by asymptotic 4-momenta of free particles of the initial $|i\rangle$ and final $|f\rangle$ asymptotic states of the collision process
@@
where $P_{i}$ and $P_{f}$ denote 4-momenta of the whole fully isolated system inside which the quantum-potential collision takes place. Having cross-sections, we can evaluate from (3) the $x$ representation of the corresponding structures, which is determined in the $p$ language. In particular, the S matrix formalism was used to estimate (in a non-relativistic approximation) the scattering cross-sections in collisions of electrons with atomic nuclei, based on an original Chyliński [43] model of internal quantum relation continuum $Q_{rel}$ space. To better understand the later let us briefly recall the singularity of Galilean space $G_{4}$. For this purpose, consider the simplest compound 2-body system $S_{M}=S_{1}+S_{2}$ configuration space $E_{3}(\mathbf{r}_{1})\times E_{3}(\mathbf{r}_{2})$ based on an arbitrary reference frame $S$ parametrizing $G_{4}$. The $G$-absolute time parametrizes evolution of all possible of $S_{M}$ generated by Hamiltonian
@@
which are plane wave states for the center of mass and bound and unbound states of the relative motion. However, these eigenstates are entangled states of the two original subsystems of $S_{M}$
@@
Due to the separability of the internal from the external degrees of freedom of $S_{M}$ it is convenient to introduce the internal relation continuum space the $\mathbf{Q}_{rel,4}^{G}$ composed of internal space $\mathbf{Q}_{rel,3}^{G}$ spanned by a relative coordinate $\mathbf{y}^{G}$ and the internal time $\tau^{G}$ which up to the translation constant, can be identified with the absolute Newtonian time $t^{G}$. It is not hard to see that the symmetry of (5) consists of rotations in $\mathbf{Q}_{rel,3}^{G}$ can be interpreted as the symmetry of empty space $\mathbf{Q}_{rel,4}^{G}$. Consequently all internal characteristic of $S_{M}$ in $\mathbf{Q}_{rel,3}^{G}$ are Galilean absolute. Note that $\mathbf{Q}_{rel,4}^{G}$ and $G_{4}$ deal with the same Galilean absolute intervals $r^{G}=|\mathbf{y}^{G}|$ and $\Delta r^{G}$, in addition the points $(\mathbf{y}^{G},\tau^{G})$ of $\mathbf{Q}_{rel,4}^{G}$ are determined through the point (events) of $G_{4}$. Thus the G-symmetry admits a relational absolute space $\mathbf{Q}_{rel,4}^{G}$ which is the Cartesian product $\mathbf{Q}_{rel,4}^{G}=\mathbf{Q}_{rel,3}\times T$, which operates with two absolute intervals

### v3

@@ 4 Quantum Information Puzzle
Although so far, there is no formal definition of quantum information, the mathematical structure of Hilbert space imposes strong constraints on processing of quantum information [70]: There is no universal broadcasting machine capable of broadcasting arbitrary quantum states. Note that in contrast to quantum information, its mathematical representation $\psi$ can be broadcast. Thus, we would say that quantum information does exist, yet it is not just the wave function, but it is represented by it. Below, we present the examples that support this view against the narrow Copenhagen treatment of the wave function (from [69]). A paradigmatic example is quantum cryptography. Namely, the key distribution schemes show that there is a feature that is naturally ascribed to the quantum information carriers (QIC). It is clear that there must be quantum communication between Alice and Bob to achieve secret key [57]. Indeed, Alice cannot simply send a wave function $\psi$ written on paper to Bob, as Eve could read it out without disturbing it. However, by sending QIC, Alice and Bob can achieve something impossible within the “classical” world. Another example is the direct calculation of quantum state function $f(\varrho)$ [63]. It can be viewed as a kind of quantum computing operating on purely quantum input. This approach is natural if the quantum state $\varrho$ is viewed as representing real quantity to be processed rather than state of our knowledge.
The above examples show clearly that the quantum information exists physically, and it cannot be identified with its mathematical representation $\psi$. Curiously, quantum entanglement - a special form of quantum information is relative due to the relativity of the concept of physical system [71, 72, 73, 74] and it is a specific feature of the composite system structure [75, 76, 73]. It depends on the choice of factorization of the underlying composite Hilbert space. For example, if we decompose the Hilbert space $H$ of the composite system $S$ as a tensor product of smaller spaces $H=H_{1}\otimes H_{2},\ldots,\otimes H_{N}$ then the unit vectors and the states of the system can be very different, so different decompositions can have different properties. This difference is due to the fact that the symmetry group $U(d_{1})\times\ldots\times U(d_{k})$, $d_{i}=\dim(H_{i})$ is smaller than $U(d)$, $d=\prod^{k}_{i=1}d_{i}$. In connection with this, the concept of quantum reference frame was proposed according to which “a particular division of a composite system into subsystems constitutes a kind of reference frame called a meronomic reference frame [77]. In particular, it has been shown that in one quantum meronomic frame, separable (product) states may become entangled and vice versa [77, 78].
In should be noted here that the relative character of quantum entanglement appears also when quantum mechanics is combined with special relativity. In particular recently it has be demonstrated that the existence of the entanglement encoded in the wave function is not universally agreed upon by different moving observers [79, 80]. This, in particular, leads to ambiguity in interpreting the entanglement as a physical resource and in connection with this it was suggested that a true resource should be quantified in a basis-independent way [79].
Note that in the case for quantized electromagnetic fields, the “quantum” and “classical” features of quantum information depend on the preparation of the initial states and the method of detection [81]. Namely, in contrast to classical communication, where the information is encoded in classical properties of electromagnetic field (intensity or phase) quantum communication requires to send quantum information carriers, such as nonclassical states of light, e.g., squeezed states or Fock states that carry a well-defined number of photons. In the typical communications links, the task is an optimization overall ensembles of input quantum states used to carry information under relevant physical constraints. It is the Holevo theorem that provides a tight bound on the mutual information attainable for a given ensemble of input quantum states.
5 Laboratory Space Versus Configuration-Space and Local Realism
@@
It still remains unclear why the NRQ model predicts the existence of stable quantum structures and EPR correlations and, at the same time, is consistent with RQF. Note that to protect the latter from superluminal speeds, the a priori postulate of local commutativity is necessary [82]. However, the RQF is not strictly limited by the speed of light [83]. In this context, the existence of the EPR correlations naturally lead to the question of their physical nature and still is subject to a wide debate [84, 29]. Interestingly, Gisin suggested “that quantum correlations somehow arise from outside space-time”[85]. Clauser and Shimony had previously suggested that “Either we must completely abandon the realistic philosophy of almost all practicing physicists or fundamentally revise our concept of space-time” [86]. The second part of the above alternative is a big challenge since Lorenz’s symmetry imposed by heavy devices is an a priori measurement condition of any physical model [87]. On the other hand, this challenge encounters an intriguing problem of incompatibility formulations of quantum mechanics in laboratory space and configuration-space, which has recently been the subject of deep consideration in the context of local realism and correlation EPR [88]. Here, laboratory (lab) space “is the three-dimensional space in which we live, and the space in the Euclidean geometry is understood”[88]. Every point $\mathbf{r}_{lab}$ within lab has a unique position in what is called lab space.
@@
According to Born’s original formulation, the wave function $\psi_{lab}(\mathbf{r}_{lab},t)$ represents a field that propagates in lab space and correspondingly the real-valued probability density $|\psi_{lab}(\mathbf{r}_{lab},t)|^{2}$ also propagates as a classical field. Thus, Born’s modification of the interpretation of the wave function is expressed only in its statistical interpretation. Moreover, lab-space formulation applies only to the single particle systems, and there is no strict method to extend this formulation to $n\geq 2$ particle systems. In this context, Clauser notes “This impossibility might be expected since Local Realism’s formulation and Bell’s Theorem forbids it. Correspondingly, experiments that refute Local Realism also refute a lab space formulation of quantum mechanics.”
@@
Von Neumann introduced the abstract configuration-space formulation of quantum mechanics for the system consisting of $N$ particles, which uses a very general abstract $k$-dimensional argument space for the wave function ($k=3N$), which enables a system with $k$-degrees of freedom: $\mathbf{r}_{1},\ldots,\mathbf{r}_{N}$. A configuration space wave function is defined as a complex function defined in the configuration space $\mathbb{R}^{3N}$.
$\psi_{conf,N}:\mathbb{R}^{3N}=\mathbb{R}^{3}\times\cdots\times\mathbb{R}^{3}\to\mathbb{C},\quad\psi_{conf,N}(\vec{\mathbf{r}}_{1},\ldots,\vec{\mathbf{r}}_{N})\in\mathbb{C}$
Clauser argues that the quantum configuration space formalism far surpasses the lab space formalism which is limited to describing only a single particle system and he concludes that quantum mechanics requires an explicitly unambiguous configuration-space formulation. On the other hand, he notices: “if we assume, according to the Copenhagen Interpretation of quantum mechanics that abstract configuration-space is inherently impenetrable, then it allows for magical action at a distance, and no objects are necessary to exist as things in the laboratory space” [88].
@@ 6 Quantum Relational Space Hypothesis
In this section, we introduce the hypothesis of existence of quantum relational space (QRS). We start with the observation that, in accordance with the correspondence interpretation of the wave function we can claim that the configuration wave function $\psi_{conf}$ is a mathematical representation (image) of quantum information. Consequently, the latter must have a non-spatial nature. Unfortunately the $\psi_{conf}$ space formalism itself does not provide any conceptual model to understand its “inner workings”, although intuitively, it seems to hide a deeper level of reality. Despite the vast knowledge about entanglement, it is not still clear where it has its habitat and how it “coexists” with fundamental Galilean and Lorentz-Poincare symmetries. In 1999, Gisin’s team conducted a Michelson-Morley-like experiment which showed that the 2-photon interferences were still visible, independently of the relative velocity between Alice’s and Bob’s reference frames (actually, the magnitude of the velocity was constant at $100m/s$, but its orientation varied) [89].
Gisin’s experiment suggests that the von Neumann’s collapse of the quantized electromagnetic field cannot be viewed as a real causal process and quantization of the field definitely renders it no longer describable as “real stuff in real space-time”, [88, 90]. In a result, Clauser concludes that ”entanglement is not allowed in lab space” [90]. While he points out that the quantum configuration-space formalism captures $N$-particle entanglement. Intuitively, the multipartite nature of entanglement implies the existence of associated internal relational structure [91].
Note that the notion of relation is more general than the notion of interaction. Remarkably, the entanglement structure encodes quantum relations that can exist even when there are no interactions, which implies non-local Bell violating correlations. It suggests that the nature of quantum information is, in principle, relational. Therefore, it seems reasonable to adopt the QRS hypothesis:
{mdframed}
The fundamental physical interactions between quantum objects generate directly unobservable relations, which precede quantum correlations, and are the source of physical properties of quantum objects.
@@
Here we assume that every completely isolated system can be associated with an intrinsic quantum space $Q_{rel}$ of directly unobservable relations generated by unitary interactions. We assume that the space of quantum relations involves intrinsic quantum degrees of freedom such as spin, color, and smell, and it precedes the space of directly observable correlations $Q_{cor}$, which is some set of events of the Minkowski $M_{4}$ space-time (Fig.1).
@@
Figure 1: $Q_{rel}$ relation space of “pre-measurements” precedes the space of correlation events $Q_{cor}$. Arrows indicate irreversible decoherence
We are not able to determine the mathematical structure of the hypothetical relational space $Q_{rel}$. The simplest way to get around this difficulty is to assume that the relational space structure is isomorphic to the Hilbert space tensorial structure. Then, we can say that quantum models provide instructions on how to model a structure of $Q_{rel}$ in terms of Hilbert space formalism. In the multipartite dynamic scenario, building of quantum relations corresponds to the situation when the interaction (entangler) between systems belonging to partial Hilbert spaces $H_{1}$, $H_{2}$, $\ldots$, $H_{N}$, builds the total Hilbert space $H=H_{1}\otimes H_{2}\otimes\ldots\otimes H_{N}$ of the final entangled state. Recently, the experimental implementation of forming of entanglement between the ion and the photoelectron in a helium atom was demonstrated [50]. Quite lately, by the numerical solution ab initio of the Schrödinger equation, it has been shown that building of directly unobservable quantum relations between light-field driven electronic dynamics in residual ion takes place in a short (attosecond) time before the photoelectron leaves the atom [51]. In particular, it have been shown that the delay of the photoelectrons carriers information on the entangled light-field dressed electronic dynamics in the residual ion.
Note that the algorithmic generation of genuine multipartite entangled states by coherent unitary deformation operations can be viewed as a kind of modeling of the $Q_{rel}$ space [92]. In particular, the pure state $\psi$ of a composite quantum system encodes information about all relations (amplitude and phase) between its subsystems, which may be related to “the correlations without correlata” [8]. Perhaps the most adequate representation of $Q_{rel}$ space is von Neumann’s reversible pre-measurement scheme modeled by the unitary evolution coupling the measurement apparatus and the measurement system which involves information transfer caused by interaction. This interaction can be realized, in particular, by specific unitary Ozawa’s Hamiltonian evolution, which does not need a description of the measurement-induced state collapse, while it provides a cut between the measurement apparatus and the measured system [93].
Note, however, that the Hilbertian “isomorphic constraint” imposed on the of $Q_{rel}$ may be too restrictive. In particular, it goes beyond Tsirelson’s problem [94] and may rule out the description of potential “exotic” effects such as quantum gravity or “stronger nonlocal correlations“ [94]. This may require more sophisticated mathematical structures such as non-separable Hilbert space [95], quantum relations formalism [96, 94].
Interestingly, the model of quantum relational continuum space based on the configuration-space of relative coordinates and momenta, whose shapes are related to each other through unitary transformations, was proposed [39] (see Sec. 7). Note that there is no reason to assume that the internal symmetry of the $Q_{rel}$ space must be the Lorentz symmetry. Obviously, the assumption on the perfect isolation of the system is a realistic idealization, as the unitary interactions do not yield outcomes [97], thus the system cannot “recognize” itself. Hence, a full description of the pre-measurement does not imply a description of the output production. It is intimately connected with the measurement problem – “Achilles’ heel” of quantum modeling of the world [25, 26]. For additive conserved quantities, the model of the measurement was proposed based on dilated Hilbert space that includes both the degrees of freedom of the measurement apparatus and the measured system [98, 99]. In realistic scenarios, however, the observer is not able to control all the degrees of freedom environment of the system and broadcast of the results [100]. Thus, he is forced to use a formalism that includes open systems and irreversible decoherence (i.e., [101, 102]). In contrast to the view that every interaction is a measurement, [7] (see of [103, 104]) we distinguish here a class of projective, strong, and instantaneous measurements that are subject to Lorentz’s symmetry imposed by heavy bases as an a priori condition for the measurement of any physical model [87]. Such measurements will lead to an actualization of quantum relations, generating the correlation space $Q_{cor}$, which can be viewed as a set of events of the Minkowski $M_{4}$ space-time.
@@
In this sense the relation space $Q_{rel}$ precedes $Q_{cor}$ a set of the actualizations (clicks) appearing as a result of irreversible measurements with specially prepared macro-objects (detectors) in finite regions $C$ of $M_{4}$. This is compatible with the dynamic conception of Minkowski space-time, which involves irreversible processes [105] Then, it is reasonable to introduce measure $\mu(C)$ as the probability of a ‘click’ provided that a suitable detector operates in region C [106]. This seems consistent with the suggestion [107] that the quantum probabilistic structure is “overbuilt“ over $M_{4}$, while the latter is only a “carrier“ for the probabilistic measures.
It should be noted that the QRS hypothesis is compatible with Reichenbach common cause principle [108] and Gisin’s conjecture that the “quantum correlations arise from outside space-time“. Indeed, according to the QRS hypothesis, the quantum correlations have their origins in the $Q_{rel}$ space. Interestingly recent results - experimental [50] and theoretical [51] prove that quantum relations associated with the forming of the corresponding wave-packets do not emerge instantaneously, but require a short, finite time.
Finally, if the QRS hypothesis is correct, it will shed new light on what is termed in the EPR folklore as ”spooky correlations”. Namely, in correlation measurements, the original quantum relations are destroyed by strong measurements interactions with the apparatus and at the same time are actualized as quantum correlations that ”remember” their initial relational non-spatial feature of non-spacial origin. This peculiar feature makes quantum correlations to reveal as instantaneous in classical space-time. In this sense they are over cross-system and therefore they can be interpreted as observer-independent resource for quantum tasks (see of [79]).
@@ 7 Chyliński’s Quantum Relational Continuum Space
In this section, we present Chyliński’s model of a relational continuum space based on the configuration-space of relative coordinates. The NRQ model ($01/c=$ 0, $\hbar\neq 0$) maintains the classical continuum of directly observable events $X^{\mu}=(\mathbf{X},X_{0})$, $\mu=1,2,3,4$, as an elementary background of most physical models. On the other hand, according to the uncertainty relations, the exact determination of 4-momentum $P^{\mu}=(\mathbf{P},P_{0})$, $P^{\mu}(\Delta P^{c}\to 0)$ rules out the $x$ localization of its quantum carrier in $M_{4}$ since $\Delta X^{\mu}\sim h/\Delta P^{\mu}\to\infty$. It destroys in an uncontrollable way the “quantum-potential” structure $\psi$ of the isolated system under measurement. As one knows, the direct p-measurements of the scattered particles in the asymptotic region do not disturb the structure of $\psi$ state, which is consistent with the state of isolation of the investigated system. The typical momentum measurements require measuring position and inferring momentum, which, however, do not violate the uncertainty principle. Note that there are also momentum measurements that do not rely on measuring position based on the Doppler effect, the energy-sensitive single-photon detector[109].
It is not obvious that Lorentz symmetry of “heavy bases” valid micro-states when it is extrapolated to the interaction region. As we know the quantum $p$-$x$ duality establishes one-to-one correspondence between the p and x representations of a Hilbert vector $|G\rangle$ as [110]
@@
Note, however, that in this abstract formulation of the state $G$ of system $S_{M}$ the physical meaning of the $x$ and $p$ variables remain undetermined. According to the $M_{4}$ framework, the $p$’s denote four-momenta of the corresponding degree of freedom of $S_{M}$ represented in some reference frame $S$. However, 4-symmetry revealed the experimental role of the asymptotic language $p$ of relativistic kinematics, which was enhanced by relativistic $S$-matrix formalism. Thus, we can first measure the $S$ matrix elements (cross-sections) parametrized by Mandelsztam $L$-invariant variable $s_{k}$ determined by asymptotic 4-momenta of free particles of the initial $|i\rangle$ and final $|f\rangle$ asymptotic states of the collision process
@@
where $P_{i}$ and $P_{f}$ denote 4-momenta of the whole fully isolated system inside which the quantum-potential collision takes place. Having cross-sections, we can evaluate from Eq. 2 the $x$ representation of the corresponding structures, which is determined in the $p$ language. In particular, the $S$-matrix formalism was used to estimate (in a non-relativistic approximation) the scattering cross-sections in collisions of electrons with atomic nuclei, based on an original Chyliński [39] model of internal quantum relation continuum $Q_{rel}$ space. To better understand the latter let us briefly recall the singularity of Galilean space $G_{4}$. For this purpose, consider the simplest composite 2-body system $S_{M}=S_{1}+S_{2}$ configuration-space $E_{3}(\mathbf{r}_{1})\times E_{3}(\mathbf{r}_{2})$ based on an arbitrary reference frame $S$ parametrizing $G_{4}$. The $G$-absolute time parametrizes evolution of all possible of $S_{M}$ generated by Hamiltonian
@@
which are plane wave states for the center of mass and bound and unbound states of the relative motion. However, these eigenstates are entangled states of the two original subsystems of $S_{M}$ (see in this context [73]).
@@
Due to the separability of the internal from the external degrees of freedom of $S_{M}$ it is convenient to introduce the internal relation continuum space the $\mathbf{Q}_{rel,4}^{G}$ composed of internal space $\mathbf{Q}_{rel,3}^{G}$ spanned by a relative coordinate $\mathbf{y}^{G}$ and the internal time $\tau^{G}$, which, up to the translation constant, can be identified with the absolute Newtonian time $t^{G}$. Note that the symmetry of Eq. 8 consisting of rotations in $\mathbf{Q}_{rel,3}^{G}$ and translation in $\tau^{G}$ can be viewed as the symmetry of empty space $\mathbf{Q}_{rel,4}^{G}$. Consequently all internal characteristic of $S_{M}$ in $\mathbf{Q}_{rel,3}^{G}$ are Galilean absolute. Moreover $\mathbf{Q}_{rel,4}^{G}$ and $G_{4}$ deal with the same Galilean absolute intervals $r^{G}=|\mathbf{y}^{G}|$ and $\Delta r^{G}$, in addition the points $(\mathbf{y}^{G},\tau^{G})$ of $\mathbf{Q}_{rel,4}^{G}$ are determined through the point (events) of $G_{4}$. Thus the G-symmetry admits a relational absolute space $\mathbf{Q}_{rel,4}^{G}$ which is the Cartesian product $\mathbf{Q}_{rel,4}^{G}=\mathbf{Q}_{rel,3}\times T$, which operates with two absolute intervals